OpenStudy (anonymous):

Find f(10) for the exponential equation f(x)=3.5(1.10^x) Please and thank you!

OpenStudy (anonymous):

f(10)=3.5(1.10^10)

OpenStudy (anonymous):

9.07809861

OpenStudy (anonymous):

The f(x) is millions of people and the other x is years since 2012

OpenStudy (jdoe0001):

that just means, the "other x" is the INPUT, and the f(x) is the result of the input, that is the OUTPUT

OpenStudy (jdoe0001):

\(\bf f(x)=3.5(1.10^{\color{red}{ x}})\qquad x=10\qquad f(x)=3.5(1.10^{\color{red}{ 10}})\)

OpenStudy (anonymous):

Oh, okay. That makes sense. Thank you!